### v1

# Daniel Calegari and Andrea Delgado

Abstract
[Abstract]Educational Process Mining (EPM) is a data analysis technique that is used to improve educational processes. It is based on Process Mining (PM), which involves gathering records (logs) of events to discover process models and analyze the data from a process-centric perspective. One specific application of EPM is curriculum mining, which focuses on understanding the learning program students follow to achieve educational goals. This is important for institutional curriculum decision-making and quality improvement. Therefore, academic institutions can benefit from organizing the existing techniques, capabilities, and limitations. We conducted a systematic literature review to identify works on applying PM to curricular analysis and provide insights for further research. From the analysis of 22 primary studies, we found that results can be classified into five categories concerning the objectives they pursue: the discovery of educational trajectories, the identification of deviations in the observed behavior of students, the analysis of bottlenecks, the analysis of stopout and dropout problems, and the generation of recommendation. Moreover, we identified some open challenges and opportunities, such as standardizing for replicating studies to perform cross-university curricular analysis and strengthening the connection between PM and data mining for improving curricular analysis.
keywords:
Educational Process Mining, process mining, curricular analysis, data mining
articletype: Systematic Review
1 Introduction
Educational Data Mining (EDM) 15, 33 is a field that deals with the analysis of data from educational information systems such as Learning Management Systems (LMS) and administrative systems. Its primary objective is to answer critical academic questions and enhance educational practices. To achieve this, EDM employs different data mining techniques to gain insights into the educational context 33, e.g., grouping similar students based on their learning and interaction patterns and detecting students with difficulties or irregular learning processes, and also to intervene in such context, e.g., predicting an outcome and recommend a more adequate strategy.
This data-centric view of educational data is refined by the process-centric view provided by Educational Process Mining (EPM) 12. In general terms, Process Mining (PM) 42 allows analyzing the records (logs) of events associated with the execution of educational processes, as depicted in Figure 1. The analysis involves three types of activities: discovery, which uses event logs to create a process model from the causal dependencies between events; conformance, which implies verifying the correspondence of the enacted business processes concerning the expected one (through a reference model); and enhancement, which uses additional information to improve the one provided by the process model, obtaining measures such as the duration of the processes, bottlenecks, or the underuse of resources, among others.
Event logs are extracted from educational information systems and are traditionally organized according to the following principles 42:
  • A process comprises cases (traces) representing a process instance execution from start to finish. Each case has a case ID, allowing one to identify a case among others uniquely. As an example, depicted in Figure 1, it could be possible to consider as a case the approval history of a student, with the student ID identifying each case.
  • A case comprises events, i.e., an action performed in the process. Each event is associated with precisely one case. In the example, an event is the approval of a course.
  • Events can have activity, time, cost, and resource attributes. At least one attribute must be present that represents the activity carried out, i.e., its activity name. In the example, the course name can be taken as the activity name, and the approval date, student, and credits earned are additional attributes of such an event.
  • Events within a case are ordered (usually by a timestamp), determining a partial order between events of the same case. In the example, the approval date defines an ordering of the events.
Figure 1: EPM schema (from 7) and structure of an event log (both adapted from 42)
Depending on the definition of a case and its events, there are different perspectives from which the data can be analyzed. As synthesized in previous systematic literature reviews 7, 16, 40, EPM pursues different objectives, e.g., to gain a better understanding of the underlying educational process, to detect collaborative or self-regulated student’s learning behavior and difficulties, and to generate recommendations or advice for students and teachers, among others. Many of these objectives relate to analyzing learner’s behavior in the context of a course, e.g., gathering data from Learning Management Systems (LMS) usage based on actual learner’s activity recorded in the event logs 45. However, stakeholders such as study program coordinators and advisors must consider the institutional viewpoint for curriculum understanding and improvement.
An academic curriculum defines a specific learning program 29. It constrains how students are expected to take the courses 47, e.g., defining that a particular course is compulsory or must be taken before or after another. Depending on the flexibility of the curriculum, students have more or less freedom to attend courses and take exams in different order. Curricular improvements must consider students’ behaviors while trying to achieve target educational goals, such as a low dropout rate and a high percentage of students graduating on time 31.
Curriculum mining/analytics 18 focuses on collecting, analyzing, and visualizing the program- and course-level data, such as program structure and course grading, to support curriculum decision-making and quality improvement specifically. It allows analyzing the strengths and weaknesses of a curriculum by identifying its components, relations, and constraints, evaluating how they fit together, identifying potential and actual problems (e.g., blind spots, bottlenecks, and deviations of students’ trajectories), checking underlying beliefs and assumptions, and determining whether the goals have been met.
Curriculum mining is identified as an application domain of EPM 16. There are general reviews on EDM 15, EPM 7, 16, 40, and PM in LMS 45 that mentioned curricular mining. However, although it is a context of utmost importance for decision-making, no specific literature review exists. It prevents us from having an overview of the existing works (the existing reviews are not comprehensive), the challenges that arise for curricular analysis, and the existing research opportunities.
In this work, we conduct a Systematic Literature Review (SLR) following the methodological guidelines in 21, 22 to identify works on applying PM to curricular analysis. We also classify and analyze the existing work to evaluate the PM scope in this domain and provide insights for further research. In particular, we briefly discuss EDM initiatives that can be connected to improve curricular analysis.
The rest of the paper is organized as follows. In Section 2, we describe the conducted literature review methodology, and in Section 3, we synthesize the findings. Then, in Section 4, we discuss the findings and investigate future research directions. Finally, in Section 6, we provide the conclusions of this work.
2 Literature review methodology
We conducted a Systematic Literature Review (SLR) based on 21, 22. As depicted in Figure 2, it defined three phases: planning, conducting, and reporting the review, the first two divided into several steps that are summarized next.
Figure 2: SLR research methodology summary
2.1 Planning the review
The need for a systematic review is identified in this phase, and the protocol for conducting the review is defined. The protocol includes the formulation of the research question, the definition of the search terms and string, the search sources to be used, and the inclusion and exclusion criteria that will be applied to identify the primary studies of the review.
In this work, the need for a systematic review is identified in the context of EDM and EPM initiatives we were carrying out within the Computer Science career of the Instituto de Computación (Computer Science Institute, InCo) of the Facultad de Ingeniería (Engineering School, FING) of the Universidad de la República (UdelaR), to be used as the basis for the career analysis and students paths towards the defined curricula. The result of this review allows us to identify critical elements to be considered in such analysis. It also made us realize that the lack of standardization between studies makes it very hard to compare curricula of the same career from different countries, even from other universities of the same country.
This work focuses on EPM initiatives, for which we defined the review protocol presented next.
Define research question The general research question guiding the review is
{mdframed}
What are the research papers on applying PM to curricular analysis (i.e., EPM papers) ?
Since it involves an understanding of the objectives pursued by related articles and the techniques they used, as well as identifying possible open research opportunities, we derived specific questions to drive the analysis:
What are the research objectives of these works?
How PM techniques are used?
What research methodologies and PM methods and tools are used?
What are the open challenges and research opportunities?
Search keywords, strings, and sources We defined a search string for the selection of published papers. We search for keywords in existing literature reviews 16, 7 and in some papers referred by them. Since many heterogeneous keywords arise, e.g., curriculum, dropout, bottleneck, registration, and trajectories, we used the general string
“process mining” AND “education*”
which considers education, educational, and education data. These keywords lead to more false positive matches, but they help cover a broader spectrum of the literature. False positives are filtered out using the inclusion and exclusion criteria introduced in the next step and the manual inspection of articles. We identified IEEE Xplore, ACM Digital Library, SpringerLink, ScienceDirect, Wiley, Scopus, and Web of Science as sources.
Inclusion and exclusion criteria The inclusion criteria defined are: (1) the paper concerns PM to curricular analysis, and (2) it is written in English. The exclusion criteria are as follows: (1) the paper is not electronically available on the web, and (2) the paper refers to some non-peer-reviewed publications, such as technical reports, book chapters, proceedings’ prefaces, and journal editorials. We did not exclude short papers since we were interested in extracting exciting ideas beyond the availability of tools or the detailed implementation of case studies.
2.2 Conducting the review
Based on the defined protocol, we conducted the review following the steps already summarized in Figure 2. In what follows, we describe the first four steps. The last step concerning the data analysis is thoughtfully described in Section 3 and Section 4.
Execute the searches The initial search returned 3.251 articles, covering papers until 2th November 2023, as shown in Table 1.
Apply inclusion & exclusion criteria We applied inclusion and exclusion criteria to the obtained papers by reading the title, abstract, and keywords. This refinement gives a primary selection of 80 papers (with duplicates), as shown in Table 1.
Select primary studies We selected the primary studies based on reading the full text of the paper. We reinforced the results by backward searching references, searching for authors’ and conferences’ web pages, Google Scholar, and the DBLP database. It allowed finding other papers marked as “Other” in Table 1, giving 34 papers. After removing duplicates, we finally identified 22 primary studies. In the last column of Table 1, we keep one of the sources for duplicates (Scopus and Web of Science are the sources that have duplicates already in the other sources). Primary studies add the specific curricula analysis perspective to the three existing literature reviews that partially refer to the topic 7, 16, 40.
Extract data We extracted the primary data of each work summarized in Table 2 (ordered by year/author). We also extracted data that allowed us to categorize the works and answer the research questions that will be analyzed in the following sections [^1] List of relevant studies, primary selection, and extracted attributes are provided as supplementary material..
Table 1: Data sources and results for literature search.
Source Total results Relevant studies Primary studies w/duplicates Primary studies
IEEE Xplore 41 17 5 5
ACM Digital Library 339 1 0 0
SpringerLink 2331 10 3 3
ScienceDirect 14 2 0 0
Wiley 147 7 0 0
Scopus 242 33 11 7
Web of Science 137 10 8 0
Other 7 7
Total 3251 80 34 22
Table 2: Primary studies
Ref Year Main author Title Source
41 2009 Trcka N. From local patterns to global models: Towards domain driven educational process mining IEEE
2 2012 Anuwatvisit S. Bottleneck mining and Petri net simulation in education situations IEEE
4 2012 Ayutaya N. Heuristic mining: Adaptive process simplification in education IEEE
29 2012 Pechenizkiy M. CurriM: Curriculum mining Scopus
10 2014 Cairns A. Towards Custom-Designed Professional Training Contents and Curriculums through Educational Process Mining Other
11 2014 Cairns A. Using semantic lifting for improving educational process models discovery and analysis Scopus
6 2015 Bendatu L. Sequence Matching Analysis for Curriculum Development Other
47 2015 Wang R. Discovering process in curriculum data to provide recommendation Other
13 2017 Cameranesi M. Students’ careers analysis: A process mining approach Scopus
39 2017 Schulte J. Large scale predictive process mining and analytics of university degree course data Scopus
9 2018 Caballero J. Discovering bottlenecks in a computer science degree through process mining techniques IEEE
48 2018 Wang R. Sequence-based Approaches to Course Recommender Systems Springer
35 2019 Salazar J. Describing educational trajectories of engineering students in individual high-failure rate courses that lead to late dropout Scopus
36 2019 Salazar J. Influence of Student Diversity on Educational Trajectories in Engineering High-Failure Rate Courses that Lead to Late Dropout IEEE
27 2021 Martinez P. Modelling Computer Engineering Student Trajectories with Process Mining Other
37 2021 Salazar J. Backpack process model (Bppm): A process mining approach for curricular analytics Scopus
38 2021 Salazar J. Curricular analytics to characterize educational trajectories in high-failure rate courses that lead to late dropout Scopus
25 2022 Mahammed N. A System to Search and Recommend Learning Courses Sequences Other
46 2022 Wagner M. A Combined Approach of Process Mining and Rule-based AI for Study Planning and Monitoring in Higher Education Springer
30 2023 Potena D. Evidence-based Student Career and Performance Analysis with Process Mining: a Case study Other
19 2023 Hobeck R. Process Mining on Curriculum-Based Study Data: A Case Study at a German University Springer
32 2023 Rafied M. Extracting Rules from Event Data for Study Planning Other
3 Data synthesis & results
This section synthesizes and analyzes the collected data, answering the first two research questions. Figure 3 summarized primary data; it depicts the distribution of the research papers through the years, the countries of the affiliation of researchers, and the type of publication (conference or journal).
Figure 3: Data summary: number of papers published by year, countries of affiliation of researchers, and kind of publications
As can be seen, the topic has been studied regularly for the last ten years. The interest is spread in universities in fourteen countries from Asia, Europe, Oceania, and South and North America. Works have been published primarily at conferences. In this case, until 2015, papers were published in more general venues. Since 2015, papers have been published in more specific EDM and education venues, such as the International Conference on Educational Data Mining and the International Conference on Learning Analytics & Knowledge. Moreover, since 2022, papers have been published in a more dedicated venue: the Education meets Process Mining (EduPM) workshop at the International Conference on Process Mining (ICPM). In the case of papers published in journals, even for recent publications from 2022, it can be seen that there are published in non-specialized journals in the EPM, EDM, or the education domains, but on journals focused on more general domains, such as the Database and Expert Systems Applications, Applied Sciences, and the Turkish Journal of Computer and Mathematics Education.
(RQ1) What are the research objectives of these works?
The main objective addressed in these papers is educational trajectories discovery. It means discovering a process model that reflects the trajectory of students, considering a trajectory as the set of courses, exams, or curriculum milestones taken or approved by students on a given period. It also involves identifying the many different variants of students’ behavior and the students’ characteristics that describe such variants. The discovery of educational trajectories is strongly related to the formal modeling of an academic curriculum from predefined requirements. It considers the curriculum’s structure, prerequisites, and business rules, e.g., courses that cannot be taken in parallel, deadlines for approval, and a maximum number of courses that can be taken in one semester.
Using a discovered or pre-authored curriculum model, it is possible to perform curricular conformance checking to check whether the observed behavior of students matches their expected behavior to identify deviations. As with the discovery of trajectories, it also involves analyzing variants of students’ behavior. This objective could be used to understand how the curriculum’s structure changes when certain administrative decisions are taken, e.g., allowing students to take courses without having all the prerequisites.
Even without deviations from the expected curricular trajectory, there could be problems with students’ progress. Bottleneck analysis focuses on identifying and analyzing causes delaying progress (bottlenecks) on the student’s educational trajectories. The stopout/dropout analysis is a more specific analysis, not focused on delays, but on identifying and analyzing causes that lead to stopout and dropout of students. In both cases, the cause analysis is focused on identifying the courses/exams, students’ characteristics, and behavioral patterns that lead to such problems.
Finally, trajectory/curriculum recommendation is devoted to making personalized recommendations to the student’s educational trajectories or the curricular design to improve the expected results, e.g., to earn a degree. These recommendations are based on the information from the analysis described before.
Table 3 summarizes a categorization of the objectives followed by the primary studies. Works are present in many categories when their focus is spread.
Table 3: Papers focused on each research objective.
Objective # References
Educational trajectories discovery: discovering a process model that reflects the educational trajectory of students. 13 41 2 4
29 11 10
47 13 39
9 27
30 19
Curricular conformance checking: checking whether the observed students’ educational trajectories match their expected curricula. 10 41 4 29
11 47 6
13 46
30 19
Bottleneck analysis: identifying and analyzing causes delaying progress (bottlenecks) on the student’s educational trajectories. 5 29 2 9
27 30
Stopout/dropout analysis: identifying and analyzing causes that lead to stopout and dropout of students. 6 29 35
36 27
37 38
Trajectory/curriculum recommendation: making personalized recommendations to the student’s educational trajectories or the curricular design to improve the expected results. 6 29 10
48 25
46 32
Research objectives can take the form of questions. As analyzed in 50, whether the PM methodology proposes to start with concrete questions to answer or if there is an exploratory analysis addressing general process-related and data-related questions, a project iteratively refines and generates domain-specific questions of interest. Some works 39, 29, 11, 13, 47, 38, 30 describe curricular analysis specific questions of interest, e.g.,
  • What is the real academic curriculum (study program)?
  • Is there a typical (or the best) way to traverse the curriculum?
  • How does the students’ progression in their first year impact the final grade and graduation time?
  • Are there patterns that capture significant differences in the behavior of students who graduated vs. those who did not?
  • Are there patterns that capture significant changes in students’ behavior from one generation to the other?
  • Which are the best practices and worst practices followed by students?
  • Do current prerequisites between courses make sense?
  • Are the particular curriculum constraints obeyed?
  • Are there patterns that describe a bottleneck in the curriculum?
  • How likely is it that a student will finish the studies successfully or drop out?
  • For late dropout students, how can the educational trajectories that consider the group of high-failure rate courses the student must retake be described?
  • What is a student’s expected time to finish the degree?
  • Are students following the study program more likely to graduate on time?
  • What are the more advantageous courses to take next?
  • Should students take some trajectory to maximize their Grade Point Average?
  • Can a student expect a similar outcome next semester, or must adjust its commitment, engagement, and overall workload?
  • Is it better to drop a course to manage expectations rather than to (predictably) fail and perhaps have to leave the degree altogether?
Two observations arise from these questions. First, we can identify student-related and curriculum-related questions focused on different stakeholders. For example, “What is a student’s expected time to finish the degree?” is focused on providing the student with knowledge about how much time they have left in their career. On the other hand, “Are students following the study program more likely to graduate on time?” seems to be a similar question, but it is focused on knowing if the study plan design allows achieving the goal, assuming the students’ specific usual behavior. Second, we can notice that the answer to some questions can be directly connected to some of the objectives, e.g., curricular conformance checking is directly associated with answering if the particular curriculum constraints are obeyed. However, other questions enclose a combination of requirements that need to be addressed with a broad perspective, as will be discussed in Section 4.
(RQ2) How PM techniques are used?
In general terms, as described in 29, curriculum mining refines the general PM tasks within the educational context: building a curriculum model from the observed student’s behavior (discovery), checking whether this behavior matches the expected curricular definitions (conformance checking), and projecting information extracted from the observed data onto the model to perform further analysis and recommend further actions (extension). Table 4 summarizes the type and perspective each work addresses.
These aspects can be further improved by combining PM with data mining and Artificial Intelligence (AI) techniques as in 46, powering the analysis and providing active feedback to students and program coordinators.
As already described and summarized in Table 3, the main objective addressed in the papers is the discovery of educational trajectories. It requires describing a process model that reflects the trajectory of students. Cases are usually actual students’ trajectories in which events of the discovered process represent courses, e.g., in 47, 2, 4, 13, 39, 9, 27, 19, or certain curriculum milestones such as the end of a semester 30. However, the semantics of an event can be conceived from different perspectives, not only from a curricular standpoint, e.g., the courses a student approves, but also from a contextual perspective, e.g., graduation in time, late dropouts, or the number of pending studies. This interpretation is used in works focusing on stopout/dropout analysis, e.g., 35, as we will describe later.
Since a trajectory refers both to a student’s path and a set of students, in some cases, there is an a priori clustering of traces based on considering additional information, e.g., demographic information or students’ performance characteristics, which sometimes requires additional data mining tools. In other cases, clustering is performed beforehand during the analysis. In 10, the authors propose a clustering method for partitioning training paths based on key performance indicators (KPIs). Moreover, in 13, the authors segmented traces into three disjoint groups depending on the final grade mark and length of the study period. In both cases, the partitioning requires analyzing first such KPI and path lengths.
When logs contain low-level information unsuitable for analyzing the problem, e.g., to analyze how knowledge areas spread over several courses are covered, it is possible to associate more precise semantics to those events (semantic lifting), e.g., associating the knowledge areas to the events 44. In 11, the authors propose a (semi)automatic procedure to associate semantics to training labels based on an ontology and then discover simplified training paths using the semantic information. This semantic lifting is also used in other works, such as 35, not to discover a trajectory but to perform a stopout/dropout analysis from another perspective.
A curriculum process model can also be formally modeled from scratch. In 41, the authors propose a method based on Colored Petri Nets patterns and show how pattern mining can be used to discover a curriculum model.
Curricular conformance checking is based on traditional conformance checking, except in 6, in which the authors check the curriculum design with the actual trajectories using sequence matching analysis. As described in 46, conformance is used in two senses: first, to identify how different cohorts behave and to identify outliers, and second, to measure adherence to a planned curriculum. As described in 19, trajectories have a high variance. Thus, two strategies are followed: clustering of traces once a cohort or specific group of students is identified, as described before, and projecting into the model and the log partial information or stages in the curricula, e.g., compulsory or first-year courses.
In the case of bottleneck analysis, in all works 2, 9, 27, 30, a process model directly represents the students’ actual trajectories, i.e., each event is a course or exam taken or approved. Since time matters, it is firmly based on performance analysis. In 2, 27, the authors identify courses that prevent academic progress by performing a waiting time analysis and identifying courses that exhibit approval times above the expected. In 9, the authors obtained frequencies of students according to the time students invest in completing a course and extracted requirements that define what a bottleneck is from the observed behavior, e.g., courses with frequencies for total passed students lower than 50%. In 30, the authors detect exam-taking patterns and use prediction models to elucidate the impact of different factors on graduation times.
In the case of stopout/dropout analysis, a process model can directly represent the students’ courses taken or approved. In 27, the authors follow this idea and identify the last courses with which students graduate or drop out. They consider if two years have passed since the last approval, a rule defined by their university. However, other works use different semantics for events. In 35, the authors built an event log at the semester level, converting approval data to higher-level events such as a STOPOUT event if the student did not enroll in courses in the semester and a BAG event if the student enrolls but has pending courses, among others. They also consider final events such as EARLY-DROPOUT, for students who dropped out no later than the fourth semester, and LATE-GRAD for students who took longer to graduate than the time defined for the program, among other events. In 36, the authors add other events of interest to see how pending courses accumulate in a semester and if that affects the stopout in the long run. They also studied how gender, income, and entry math skills may explain differences in trajectories. In 37, they refine this idea by defining a model according to the courses students have failed and have yet to pass. Finally, they focus on dropout in 38, using a state diagram that shows the possible result events in the retake-stopout model.
Finally, trajectory/curriculum recommendations, such as in 29, 46, are based on combining several strategies. In general terms, as described in 46, PM techniques are used to identify successful students’ paths and to detect deviations from expected plans. This information is combined with rule-based AI strategies to derive rules as recommendations of more suitable study paths with higher success rates, i.e., predictions with a certain confidence level. Many possible strategies are applied, e.g., in 48, the authors apply sequence-based approaches (e.g., sequential pattern mining); in 25, the authors use logistic regression for resolving a classification problem; and in 32, the authors derive decision tree models. In a less restrictive curricular context, one possibility is to discover the trajectories that students naturally follow and define a curriculum based on that. In 10, the authors analyze social networks between courses and training providers of a consulting company, e.g., to discover popular providers. They also propose a clustering method for partitioning training paths based on key performance indicators and use discovery to analyze such clusters.
(RQ3) What research methodologies and PM methods and tools are used?
In Table 4, there is a summary of the PM strategies followed, indicating the type of PM: Discovery (D), Conformance checking (C), and Extension (E). The primary type is discovery since most works tend to discover student trajectories (control-flow perspective). Few works also focus on the time perspective when considering enhancement 2, 29, 9, 27, e.g., bottleneck analysis, and only one paper 10 considers the organizational perspective to provide recommendations.
As tangible results of the works, most of them provide real-world application examples by taking historical information from universities and consulting companies supporting software systems. From a methodological perspective, since 2019, some works have started using the PM2 43 methodology to guide the execution of their PM projects; before that, only a few works described using an ad hoc methodology.
Only some works define comprehensive tools supporting curricular mining initiatives. In 29, the authors envision CurriM, a tool covering many PM actions about curricular models. Moreover, in 46, the authors describe an approach and tool to support individualized study planning and monitoring using campus management system data and study program models. In this case, the StudyBuddy tool provides a planning tool for students, while the BuddyAnalytics provides an analytics dashboard for program designers. Unfortunately, none of these tools are available. However, the project described in 46 is ongoing. Other works also refer to unavailable developed tools supporting their specific analysis.
Before 2017, the most used tool for performing analysis was ProM. Some works describe the discovery algorithms used, i.e., Fuzzy Miner, 11, Heuristic Miner 11, 27 and infrequent Inductive Miner (iIM) 13. Moreover, in 4, the authors analyze the appropriateness of the Heuristic Miner for process discovery. Since 2017, there has been a slight change in the appearance of other PM tools, i.e., Disco and bupaR. Also, in 39, the authors describe the combination of PM and DM tools they used: KNIME, Disco, ProM, and R. Finally, in 47, the authors developed a curriculum simulator to mimic the behaviors of different kinds of students, which is unavailable.
Table 4: PM strategies.
Paper Year Result Tool Method PM Type Perspective
41 2009 example, tool ProM D, C control-flow
2 2012 example ProM Ad hoc D, C, E control-flow, time
4 2012 example ProM Ad hoc D control-flow
29 2012 tool ProM D, C, E control-flow, time
10 2014 example, tool ProM D, E control-flow, org.
11 2014 example ProM D, C control-flow
6 2015 example, tool ProM Ad hoc C control-flow
47 2015 tool D, C control-flow
13 2017 example ProM D, C control-flow
39 2017 example, tool ProM, Disco D control-flow
9 2018 example ProM Ad hoc D, E control-flow, time
48 2018 example, tool D control-flow
35 2019 example bupaR PM2 D control-flow
36 2019 example bupaR PM2 D control-flow
27 2021 example ProM D, E control-flow, time
37 2021 example bupaR PM2 D control-flow
38 2021 example bupaR PM2 D control-flow
25 2022 tool D control-flow
46 2022 example, tool D, C control-flow
30 2023 example Pm4Py D, C control-flow
19 2023 example ProM, Disco PM2 D, C control-flow
32 2023 example, tool Pm4Py D control-flow
4 Opportunities & challenges
In this section, we discuss (RQ4) What are the open challenges and research opportunities? Since curricular analysis is within a broader context, there are general PM challenges from a technical 42 and organizational 26 perspective, as well as more specific EPM challenges 7 that must be addressed. In what follows, we summarize and complement specific curricular analysis challenges described in primary studies.
Heterogeneity and complexity.
PM tools must handle voluminous data from educational information systems. This data describes heterogeneous and complex situations since educational processes reflect the high diversity of behaviors in students’ learning paths 10. Additionally, some undesired behaviors from the PM perspective are not extraordinary, e.g., non-finalized cases of dropout students. Moreover, as described in 19, curricula are often flexible, particularly when there is a wide variety of compulsory courses and when exceptions are considered, e.g., exams could be repeated after failing, or there are lateral program entries when a student switches its career or spend a semester studying in a foreign program. All these aspects affect process discovery (generating spaghetti models) and conformance checking. Several strategies are followed, such as identifying consistent student cohorts or specific groups of students (clustering) with homogeneous characteristics and filtering them, reducing the complexity of the discovered process models.
Cluster analysis aims to identify typical behavior patterns of groups of similar students based on intra- and inter-course data such as examination results and order of the courses in their trajectories 24, 23. In 11, 13, the authors mentioned the need to develop new and refined clustering and classification techniques and to compare results from different techniques in the context of curricular analysis.
Clustering and filtering strategies could be applied on demand within an iterative analysis, e.g., performing trace clustering on students with similar characteristics and filtering a subset of such traces by looking at other specific student attributes. Another strategy is aggregating events when the raw data might need to be more cohesive or only considering partial information or stages in the curricula, e.g., compulsory or first-year courses, to perform the analysis. As noticed in 19, the use of filters to reduce the noise of the log and the resulting models comes at the cost of reducing their expressiveness.
Integrated analysis strategies.
Some analyses, e.g., analyzing dropout, sometimes require links with other contextual aspects, such as demographic information of students and the context in which a course is offered (e.g., the number of teachers assigned to a course). In this context, two requirements arise: enriching logs with additional data and incorporating multidimensional and integrated analysis strategies between PM and data mining (DM), as devised in 14. DM allows projecting information from the observed data onto the models to understand underlying relationships 29. In this sense, it is possible not only to project simple demographic factors 17 such as gender, age, nationality, ethnicity, family education, or economic position but to complement and complete curricula analysis using the educational history of the students and vocational ambitions to explain differences in trajectories 34, and econometric models to quantify the influence of variables along time 35. DM can also enrich the analysis of university curricula and the acquired competencies, for example, adding Tracer Study (TS) data of their graduates, such as the waiting period for getting the first job, the success of graduates competing in the selection, earned salaries, and graduates’ opinions 3. Process Cubes 8 could be explored in this multidimensional context for data management and log generation.
Curricular flexibility and evolution
. Student diversity in higher education and the need to deliver students who adapt to changes has made educational institutions address these needs by designing a flexible curriculum 20. To answer if current prerequisites make sense, it is necessary to evaluate what would happen if these prerequisites were not present, something that can only be approximated based on assumptions or through the generation of simulations. Recent studies based on DM and simulation present generalized approaches to exploring the impact of policy changes on the curriculum and creating effective and individually verifiable simulations without requiring a rigid curriculum graph to evaluate changes to flexible curricula 5. Process simulation could be applied to possible trajectories based on the history of discovered trajectories. It would also be interesting to compare the results of both types of simulation to assess their potential.
Another related concept in the context of curricular evolution is concept drift 10, 29, 39, i.e., a significant change in a process over time. PM is particularly sensitive to this aspect, and more profound studies still need to be developed in the educational context.
Prediction.
DM techniques have been widely used to predict students’ performance and dropout, using demographic and educational attributes, applying techniques such as Decision Trees, Random Forest, Bayes Network, and others 1. Since PM is also used for these purposes, comparing such results and determining how they can be combined is necessary.
Metrics.
Using metrics provides objective information on which to evaluate the current state of a curriculum and base decisions regarding policy changes. It could be possible to combine the discovered trajectories with specific curriculum metrics 28. Only a few works, such as 46, have used Key Performance Indicators (KPIs) about students, study programs, or cohorts, such as the success rate of a course and exams passed in a semester. Further works could explore and evaluate the use of such information and determine its potential for integrated analysis.
Methodologies.
At a methodological level, some works use a well-defined but general methodology, e.g., PM2, for guiding a curricular analysis project. We may extend the PM2 process to conduct concrete curricular analysis activities. It would imply defining specific actions and recommending particular techniques to be applied in each step, according to predefined questions of interest for curricular analysis. Moreover, we can consider a methodological context with an analysis based on existing combined DM and PM methods, as in 14.
Tools.
Tools must be available if the curricular analysis is expected to be interpreted by the end users, e.g., CurriM 29 and StudyBuddy and BuddyAnalytics 46. Many opportunities exist, e.g., extending an existing PM tool with a specific layer of curricular analysis techniques guided by a concrete methodology.
Providing adequate visualization techniques and notations for improving the interpretation of results by the end users is also a challenge, as identified in 10, 7, 19.
There is also space for improvement in applying predictive PM to provide recommendations. It could be possible to provide students with the ability to simulate an ideal trajectory and predict the probability of success according to it.
Replication.
Several aspects make replication of studies hard, e.g., no common methodology is followed for the same guiding questions, there are many alternatives to perform the analysis, data mining and PM strategies are usually disconnected, there is heterogeneous data, and no standard log format/content is defined, e.g., which demographic attributes should be considered for a given study. Some works, such as 46, tend to define more standardized aspects, e.g., a general data model for PM initiatives independent of academic systems and programs. Given that several referred study programs have similarities, based on a standard methodology and a set of predetermined studies, it would be interesting to perform a cross-university curricular analysis to detect similarities and differences.
5 Threats to validity
In this section, we present the threats to validity according to 49: construction, internal, conclusion, and external.
As a threat to construct validity, the selection of sources or the search string may not retrieve all relevant articles. To minimize them, we selected recognized sources based on their coverage of international publications from conferences and top-level journals in the area, and we opted for a search string that maximized the inclusion of articles using process mining and education* as critical terms. Exploratory searches were carried out in the sources and discussion among the authors for prior adjustments.
Internal validity is threatened by reproducibility and bias in the selection of studies. To minimize them, the protocol was specified, including the sources, the search string, and the inclusion/exclusion criteria, and the results were discussed between the authors at each stage to prevent possible selection bias.
As a threat to external validity, the quantity and relevance of the selected studies can affect the generalization of results. To minimize this threat, the articles were selected using the inclusion/exclusion criteria defined in the protocol and discussions between the authors. Although the number of articles selected compared to those recovered may seem low, the review’s focus is specifically on EPM over curricula analysis, which narrows the area. The results presented represent the topic’s current state to our knowledge. In any case, there could be articles that need to be included that were not published or indexed at the time of the searches.
The validity of the conclusions is threatened by dependence on the point of view of the authors conducting the review. To minimize this threat, the information recovered from the analyzed articles is the one indicated in each proposal, and if there were different visions in the extraction of the identified categories, these were discussed between the authors.
6 Conclusions
We have summarized the results of a systematic literature review that identifies works on applying PM to curricular analysis. It has been a small but active research area in the last ten years, encompassing the PM community’s development.
Works span the three types of PM (discovery, conformance checking, and extension) and can be classified into five categories, from the discovery of educational trajectories to specific analysis of common problems such as stopout and dropout and the production of recommendations based on evidence.
The objectives cannot be reached using PM in isolation. Many other data mining and Artificial Intelligence (AI) techniques could be used in combination, such as clustering of traces for different student groups to perform more specific analysis, the definition of KPIs for expressing target educational goals, and using AI strategies to derive rules as recommendations for higher success rate study paths.
Despite the existing work, we have identified some research opportunities, e.g., tightening the connecting with EDM for improving curricular analysis, improving the methodological settings to perform a curricular analysis according to predefined questions of interest, and the development of openly-available tools for end users to reduce the costs of implementing a PM initiative from scratch.